# Local Asymptotic Normality for Mixed Fractional Brownian Motion under High-Frequency Observation

### v2

@@ Abstract
In this paper we will consider the LAN property for both the Hurst parameter $H>3/4$ and the variance of the fractional Brownian motion plus an independent standard Brownian motion (called mixed fractional Brownian motion) with high-frequency observation. We will first remove the $H$-score linear term and orthogonalize the remainder through two non-diagonal transformations, then we can construct the CLT for the quadratic form base on $\|\cdot\|_{\mathrm{op}}/\|\cdot\|_{F}\to 0$. At last we obtain a diagonal Gaussian LAN expansion with an explicit information matrix. Beyond the case of $H>3/4$, we also present that the $\|\cdot\|_{\mathrm{op}}/\|\cdot\|_{F}\to 0$ method is also useful for the case of $H<3/4$ and the proof will be concise compared with the Whittle translation method. We consider that this method can be applied to this type of problem, including the fractional Ornstein-Uhlenbeck model and mixed fractiona O-U process.
@@ 1 Introduction
$Y_{t}=\sigma B_{t}^{H}+W_{t},\qquad t\geq 0,$ (1)
where $B^{H}$ is a standard fractional Brownian motion with Hurst index $H\in(3/4,1)$, $W$ is an independent standard Brownian motion, and the parameter $\theta=(\sigma,H)$ is unknown. We observe discrete high-frequency increments on a growing time horizon:
@@
$X_{n,i}:=Y_{i\Delta_{n}}-Y_{(i-1)\Delta_{n}}=\sigma\Delta_{n}^{H}G_{n,i}+\Delta_{n}^{1/2}Z_{n,i},\qquad i=1,\dots,n,$ (2)
where $(G_{n,i})$ is fractional Gaussian noise (fGn) and $(Z_{n,i})$ is standard Gaussian white noise, independent of $(G_{n,i})$. In this paper we will construct the LAN property for the unknown parameter $\theta$ under this high-frequency observation.
@@
Let $\bigl(\mathbf{P}_{\theta}^{n}\bigr)_{\theta\in\Theta}$ be a sequence of statistical experiments indexed by $\theta\in\Theta\subset\mathbb{R}^{k}$. We say it is locally asymptotically normal at $\theta_{0}\in\Theta$ if there exist non-singular $k\times k$ matrices $\phi(n)=\phi(n,\theta_{0})$ with $\|\phi(n)\|\to 0$, random vectors $Z_{n,\theta_{0}}\in\mathbb{R}^{k}$, and a symmetric nonnegative definite matrix $I(\theta_{0})$ such that for every fixed $u\in\mathbb{R}^{k}$,
$\log\frac{d\mathbf{P}^{n}_{\theta_{0}+\phi(n)u}}{d\mathbf{P}^{n}_{\theta_{0}}}(X^{n})=u^{\top}Z_{n,\theta_{0}}-\frac{1}{2}u^{\top}I(\theta_{0})u+o_{\mathbb{P}_{\theta_{0}}}(1),\qquad Z_{n,\theta_{0}}\Rightarrow\mathcal{N}\bigl(0,I(\theta_{0})\bigr).$
Denote by $W_{2,k}$ the class of loss functions $\ell:\mathbb{R}^{k}\to[0,\infty)$ that are symmetric, subconvex and satisfy
@@
Assume $\bigl(\mathbf{P}_{\theta}^{n}\bigr)_{\theta\in\Theta}$ is LAN at $\theta_{0}$ with normalization matrices $\phi(n)$ and information $I(\theta_{0})$. Then for any sequence of estimators $\hat{\theta}_{n}$, any $\ell\in W_{2,k}$, and any $\delta>0$,
$\liminf_{n\to\infty}\;\sup_{\|\theta-\theta_{0}\|<\delta}\mathbf{E}_{\theta}\!\left[\ell\!\left(\phi(n)^{-1}\bigl(\hat{\theta}_{n}-\theta\bigr)\right)\right]\ \geq\ \int_{\mathbb{R}^{k}}\ell(x)\,\varphi_{I(\theta_{0})^{-1}}(x)\,dx,$
where $\varphi_{I(\theta_{0})^{-1}}$ denotes the centered Gaussian density with covariance $I(\theta_{0})^{-1}$ (restricted to the range of $I(\theta_{0})$ when $I(\theta_{0})$ is singular).
@@ 2 Main Results
For any integrable (real-valued) symbol $u$ on $\pi,\pi]$, we define $T_{n}(u)$ be the $n\times n$ Toeplitz matrix with entries
@@
then our observation $X^{n}=(X_{n,1},\dots,X_{n,n})^{\top}$ is a centered Gaussian vector with covariance
$V_{n}(\theta)=\Delta_{n}\big(I_{n}+\gamma_{n}T_{n}(H)\big),\qquad\theta=(\sigma,H),$ (5)
@@ 2.1 Exact score functions
$S_{\sigma,n}:=\partial_{\sigma}\ell_{n}$ $=\frac{\gamma_{n}}{\sigma}\Big(X_{n}^{\top}\Psi_{\sigma,n}X_{n}-\mathbb{E}[X_{n}^{\top}\Psi_{\sigma,n}X_{n}]\Big),$ (7)
$S_{H,n}:=\partial_{H}\ell_{n}$ $=\frac{\gamma_{n}}{2}\Big(X_{n}^{\top}\Psi_{H,n}X_{n}-\mathbb{E}[X_{n}^{\top}\Psi_{H,n}X_{n}]\Big),$ (8)
@@
$\log\frac{d\mathbb{P}^{n}_{\theta+hr_{n}^{-1}}}{d\mathbb{P}_{\theta}^{n}}=h^{\top}\Xi_{n}-\frac{1}{2}h^{\top}I^{\perp}h+o_{\mathbb{P}}(1),$
@@ 3.1 Some General Properties of fGn and Score Function
Let $A_{n}:=I_{n}+\gamma_{n}T_{n}(H)$ and let $Z\sim\mathcal{N}(0,I_{n})$. Then we can write
$X_{n}=\sqrt{\Delta_{n}}\,A_{n}^{1/2}Z.$
@@
$S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}\Big(Z^{\top}C_{n}Z-\mathrm{tr}(C_{n})\Big),$ (9)
@@
$R_{H,n}:=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}Z-\mathrm{tr}(D_{n})\Big),$ (11)
@@
By the exact score representation, the weighting matrix for the $H$-score corresponds to $2\ln(\Delta_{n})\,T_{n}(H)+\dot{T}_{n}(H)$ sandwiched by $A_{n}^{-1/2}$. Hence, the centered quadratic form splits into the sum of the $2\ln(\Delta_{n})$-part and the $\dot{T}_{n}(H)$-part. Using the identity $\partial_{H}\gamma_{n}=2\gamma_{n}\ln(\Delta_{n})$ and the scaling relation $S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}Q_{n}(C_{n})$, we obtain the decomposition (10) with $R_{H,n}$ as defined in (11). ∎
@@
Let $Z\sim\mathcal{N}(0,I_{n})$ and let $M_{n}$ be a sequence of real symmetric matrices. Set $Q_{n}(M_{n}):=Z^{\top}M_{n}Z-\mathrm{tr}(M_{n})$. If
@@
Diagonalize $M_{n}=U_{n}^{\top}\Lambda_{n}U_{n}$ with $U_{n}$ orthogonal and $\Lambda_{n}=\mathrm{diag}(\lambda_{n,1},\dots,\lambda_{n,n})$. Then $U_{n}Z\stackrel{{\scriptstyle d}}{{=}}Z$ and
$Q_{n}(M_{n})=\sum_{i=1}^{n}\lambda_{n,i}(Z_{i}^{2}-1),\qquad\mathrm{Var}(Q_{n}(M_{n}))=2\sum_{i=1}^{n}\lambda_{n,i}^{2}=2\left\lVert M_{n}\right\rVert_{F}^{2}.$
Let $\xi_{n,i}:=\lambda_{n,i}(Z_{i}^{2}-1)/(\sqrt{2}\left\lVert M_{n}\right\rVert_{F})$, so that $\sum_{i}\xi_{n,i}=\frac{Q_{n}(M_{n})}{\sqrt{2}\left\lVert M_{n}\right\rVert_{F}}$ and the $\xi_{n,i}$ are independent and centered. Since $Z_{i}^{2}-1$ has a finite fourth moment, there exists a constant $C$ such that $\mathbb{E}[(Z_{i}^{2}-1)^{4}]\leq C$. Thus, we can verify the Lyapunov condition:
@@
$\ell_{H}(\lambda)=c_{H}\,|2\sin(\lambda/2)|^{p+2}\Big(|\lambda|^{-2H-1}+B(\lambda,H)\Big)=c_{H}\,|2\sin(\lambda/2)|^{2H+1}\Big(|\lambda|^{-2H-1}+B(\lambda,H)\Big).$
@@
$\dot{f}_{H}(\lambda)=\partial_{H}\!\big(|2\sin(\lambda/2)|^{-p}\big)\,\ell_{H}(\lambda)+|2\sin(\lambda/2)|^{-p}\,\dot{\ell}_{H}(\lambda)=|2\sin(\lambda/2)|^{-p}\Big(-2\ln|2\sin(\lambda/2)|\,\ell_{H}(\lambda)+\dot{\ell}_{H}(\lambda)\Big),$
@@
Let $a_{n}(\lambda):=1+\gamma_{n}f_{H}(\lambda)$, $A_{n}=T_{n}(a_{n})$ which is the Toeplitz matrix driven by the function $a_{n}(\lambda)$, for $n\Delta_{n}\to\infty$ and $b,c\in\{f_{H},\dot{f}_{H}\}$, let us define $k(f_{H}):=0$ and $k(\dot{f}_{H}):=1$ which counts the number of logarithmic Fisher–Hartwig factors (present in $\dot{f}_{H}$ but not in $f_{H}(\lambda)$)hence it governs the possible $\log n$ loss in the remainder. Then
@@
$|\dot{f}_{H}(\lambda)|\leq f_{H}(\lambda)\,\bigl(2|\log|\lambda||+C\bigr)+C\,|\lambda|^{-p}\leq C\,f_{H}(\lambda)\bigl(1+|\log|\lambda||\bigr).$
@@ 3.2 Central Limit Theorem for $S_{\sigma,n}$
$J_{2}(H,\sigma)$ $:=\int_{\mathbb{R}}w(x)\Bigl(C_{H}-2\ln|x|\Bigr)^{2}\,dx.$ (30)
@@
Now we can deal with the part of the parameter $\sigma$ of the score funtion
@@
$\left(\sqrt{2}\,\frac{\gamma_{n}}{\sigma}\left\lVert C_{n}\right\rVert_{F}\right)^{2}=2\frac{\sigma^{4}\Delta_{n}^{2p}}{\sigma^{2}}\cdot\frac{n}{2\pi}\Delta_{n}^{1-2p}\Big(J_{0}+o(1)\Big)=\frac{\sigma^{2}}{\pi}\,(n\Delta_{n})\Big(J_{0}+o(1)\Big)=\frac{\sigma^{2}}{\pi}\,\mathcal{T}_{n}\Big(J_{0}+o(1)\Big).$
@@ 3.3 The CLT for the Remainder of Score function of $H$
$\Delta_{n}\,\kappa_{n}(x)b_{n}(x)^{2}$ $=\Delta_{n}^{1-2p}w(x)\Bigl(4\ln^{2}(1/\Delta_{n})+4\ln(1/\Delta_{n})\bigl(C_{H}-2\ln|x|\bigr)+\bigl(C_{H}-2\ln|x|\bigr)^{2}\Bigr)$
$\quad+o\!\big(\Delta_{n}^{1-2p}\ln^{2}(1/\Delta_{n})\big).$ (40)
@@
$R_{H,n}:=\frac{\gamma_{n}}{2}\Big(Z^{\top}D_{n}Z-\mathrm{tr}(D_{n})\Big),\qquad Z\sim\mathcal{N}(0,I_{n}),$
@@
$\gamma_{n}^{2}\left\lVert D_{n}\right\rVert_{F}^{2}=\sigma^{4}\Delta_{n}^{2p}\cdot\frac{n}{2\pi}\Delta_{n}^{1-2p}\Bigl(4\ln^{2}(1/\Delta_{n})\,J_{0}+o(\ln^{2}(1/\Delta_{n}))\Bigr)=\frac{2\sigma^{4}}{\pi}\,\mathcal{T}_{n}\,\ln^{2}(1/\Delta_{n})\Bigl(J_{0}+o(1)\Bigr),$

### v3

@@ Abstract
In this paper we will consider the LAN property for both the Hurst parameter $H>3/4$ and the variance of the fractional Brownian motion plus an independent standard Brownian motion (called mixed fractional Brownian motion) with high-frequency observation. We will first remove the $H$-score linear term and orthogonalize the remainder through two non-diagonal transformations, then we can construct the CLT for the quadratic form base on $\|\cdot\|_{\mathrm{op}}/\|\cdot\|_{F}\to 0$. At last we obtain a diagonal Gaussian LAN expansion with an explicit information matrix. Beyond the case of $H>3/4$, we also present that the $\|\cdot\|_{\mathrm{op}}/\|\cdot\|_{F}\to 0$ method is also useful for the case of $H<3/4$ and the proof will be concise compared with the Whittle translation method. We consider that this method can be applied to this type of problem, including the fractional Ornstein-Uhlenbeck model and mixed fractional O-U process.
mixed fractional Brownian motion,
keywords:
[class=MSC]
keywords:
[…]
and
@@ 1 Introduction
$Y_{t}=\sigma B_{t}^{H}+{B_{t}},\qquad t\geq 0,$ (1)
where $B^{H}$ is a standard fractional Brownian motion with Hurst index $H\in(3/4,1)$, $B$ is an independent standard Brownian motion, and the parameter $\theta=(\sigma,H)$ is unknown. We observe discrete high-frequency increments on a growing time horizon:
@@
$X_{n,i}:=Y_{i\Delta_{n}}-Y_{(i-1)\Delta_{n}}=\sigma\Delta_{n}^{H}G_{n,i}+\Delta_{n}^{1/2}{W_{n,i}},\qquad i=1,\dots,n,$ (2)
where $(G_{n,i})$ is fractional Gaussian noise (fGn) and $(W_{n,i})$ is standard Gaussian white noise, independent of $(G_{n,i})$. In this paper we will construct the LAN property for the unknown parameter $\theta$ under this high-frequency observation.
@@
Let $\bigl(\mathbb{P}_{\theta}^{n}\bigr)_{\theta\in\Theta}$ be a sequence of statistical experiments indexed by $\theta\in\Theta\subset\mathbb{R}^{k}$. We say it is locally asymptotically normal at $\theta_{0}\in\Theta$ if there exist non-singular $k\times k$ matrices $\phi(n)=\phi(n,\theta_{0})$ with $\|\phi(n)\|\to 0$, random vectors $Z_{n,\theta_{0}}\in\mathbb{R}^{k}$, and a symmetric nondegenerate matrix $I(\theta_{0})$ such that for every fixed $u\in\mathbb{R}^{k}$,
$\log\frac{d\mathbb{P}^{n}_{\theta_{0}+\phi(n)u}}{d\mathbb{P}^{n}_{\theta_{0}}}(X_{n})=u^{\top}Z_{n,\theta_{0}}-\frac{1}{2}u^{\top}I(\theta_{0})u+o_{\mathbb{P}_{\theta_{0}}}(1),\quad Z_{n,\theta_{0}}\Rightarrow\mathcal{N}\bigl(0,I(\theta_{0})\bigr).$
Denote by $W^{2,k}$ the class of loss functions $\ell:\mathbb{R}^{k}\to[0,\infty)$ that are symmetric, nonnegative quasi-convex function and satisfy
@@
Assume $\bigl(\mathbb{P}_{\theta}^{n}\bigr)_{\theta\in\Theta}$ is LAN at $\theta_{0}$ with normalization matrices $\phi(n)$ and information $I(\theta_{0})$. Then for any sequence of estimators $\hat{\theta}_{n}$, any $\ell\in W^{2,k}$, and $\delta>0$,
$\liminf_{\delta\rightarrow\infty}\liminf_{n\to\infty}\;\sup_{\|\theta-\theta_{0}\|<\delta}\mathbb{E}_{\theta}\!\left[\ell\!\left(\phi(n)^{-1}\bigl(\hat{\theta}_{n}-\theta\bigr)\right)\right]\ \geq\ \int_{\mathbb{R}^{k}}\ell(x)\,\varphi_{I(\theta_{0})^{-1}}(x)\,dx,$
where $\varphi_{I(\theta_{0})^{-1}}$ denotes the centered Gaussian density with covariance $I(\theta_{0})^{-1}$ (restricted to the range of $I(\theta_{0})$ when $I(\theta_{0})$ is nonsingular).
@@ 2 Main Results
For any integrable real-valued symbol $u$ on $\pi,\pi]$, we define $T_{n}(u)$ be the $n\times n$ Toeplitz matrix with entries
@@
then our observation ${X_{n}}=(X_{n,1},\dots,X_{n,n})^{\top}$ is a centered Gaussian vector with covariance
$V_{n}:=V_{n}(\theta)=\Delta_{n}\big(I_{n}+\gamma_{n}T_{n}(H)\big),\qquad\theta=(\sigma,H),$ (5)
@@ 2.1 Exact score functions
$S_{\sigma,n}:=\partial_{\sigma}\ell_{n}$ $=\frac{\gamma_{n}}{\sigma\Delta_{n}}\Big(X_{n}^{\top}\Psi_{\sigma,n}X_{n}-\mathbb{E}[X_{n}^{\top}\Psi_{\sigma,n}X_{n}]\Big),$ (7)
$S_{H,n}:=\partial_{H}\ell_{n}$ $=\frac{\gamma_{n}}{2\Delta_{n}}\Big(X_{n}^{\top}\Psi_{H,n}X_{n}-\mathbb{E}[X_{n}^{\top}\Psi_{H,n}X_{n}]\Big),$ (8)
@@
$\log\frac{d\mathbb{P}^{n}_{\theta+hr_{n}^{-1}}}{d\mathbb{P}_{\theta}^{n}}=h^{\top}\Xi_{n}-\frac{1}{2}h^{\top}I^{\perp}h+o_{\mathbb{P}^{n}_{\theta}}(1),$
@@ 3.1 Some General Properties of fGn and Score Function
Let $A_{n}:=I_{n}+\gamma_{n}T_{n}(H)$ and $X_{n}=\sqrt{\Delta_{n}}\,A_{n}^{1/2}Z_{n}$. Then we have
$Z_{n}\sim\mathcal{N}(0,I_{n}).$
@@
$S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}\Big({Z_{n}}^{\top}C_{n}{Z_{n}}-\mathrm{tr}(C_{n})\Big),$ (9)
@@
$R_{H,n}:=\frac{\gamma_{n}}{2}\Big(Z_{n}^{\top}D_{n}Z_{n}-\mathrm{tr}(D_{n})\Big),$ (11)
@@
Since $X_{n}\sim\mathcal{N}(0,V_{n}(\theta))$ and
$Z_{n}=\frac{1}{\sqrt{\Delta_{n}}}A_{n}^{-1/2}X_{n},$
where $A_{n}=I_{n}+\gamma_{n}T_{n}(H)$, then $Z_{n}\sim\mathcal{N}(0,I_{n})$. By the exact score representation, the weighting matrix for the $H$-score corresponds to $2\ln(\Delta_{n})\,T_{n}(H)+\dot{T}_{n}(H)$ sandwiched by $A_{n}^{-1/2}$. Hence, the centered quadratic form splits into the sum of the $2\ln(\Delta_{n})$-part and the $\dot{T}_{n}(H)$-part. Using the identity $\partial_{H}\gamma_{n}=2\gamma_{n}\ln(\Delta_{n})$ and the scaling relation $S_{\sigma,n}=\frac{\gamma_{n}}{\sigma}Q_{n}(C_{n})$, we obtain the decomposition (10) with $R_{H,n}$ as defined in (11). ∎
@@
Let $Z_{n}\sim\mathcal{N}(0,I_{n})$ and let $M_{n}$ be a sequence of real symmetric matrices. Set $Q_{n}(M_{n}):=Z_{n}^{\top}M_{n}Z_{n}-\mathrm{tr}(M_{n})$. If
@@
Diagonalize $M_{n}=U_{n}^{\top}\Lambda_{n}U_{n}$ with $U_{n}$ orthogonal and $\Lambda_{n}=\mathrm{diag}(\lambda_{n,1},\dots,\lambda_{n,n})$. Then $U_{n}Z_{n}\stackrel{{\scriptstyle d}}{{=}}Z_{n}$ and
$Q_{n}(M_{n})=\sum_{i=1}^{n}\lambda_{n,i}(Z_{n,i}^{2}-1),\qquad\mathrm{Var}(Q_{n}(M_{n}))=2\sum_{i=1}^{n}\lambda_{n,i}^{2}=2\left\lVert M_{n}\right\rVert_{F}^{2}.$
Let $\xi_{n,i}:=\lambda_{n,i}(Z_{n,i}^{2}-1)/(\sqrt{2}\left\lVert M_{n}\right\rVert_{F})$, so that $\sum_{i}\xi_{n,i}=\frac{Q_{n}(M_{n})}{\sqrt{2}\left\lVert M_{n}\right\rVert_{F}}$ and the $\xi_{n,i}$ are independent and centered. Since $Z_{n,i}^{2}-1$ has a finite fourth moment, there exists a constant $C$ such that $\mathbb{E}[(Z_{n,i}^{2}-1)^{4}]\leq C$. Thus, we can verify the Lyapunov condition:
@@
$\ell_{H}(\lambda)$ $=c_{H}\,|2\sin(\lambda/2)|^{p+2}\Big(|\lambda|^{-2H-1}+B(\lambda,H)\Big)$
$=c_{H}\,|2\sin(\lambda/2)|^{2H+1}\Big(|\lambda|^{-2H-1}+B(\lambda,H)\Big).$
@@
$\dot{f}_{H}(\lambda)$ $=\partial_{H}\!\big(|2\sin(\lambda/2)|^{-p}\big)\,\ell_{H}(\lambda)+|2\sin(\lambda/2)|^{-p}\,\dot{\ell}_{H}(\lambda)$
$=|2\sin(\lambda/2)|^{-p}\Big(-2\ln|2\sin(\lambda/2)|\,\ell_{H}(\lambda)+\dot{\ell}_{H}(\lambda)\Big),$
@@
Let
$a_{n}(\lambda):=1+\gamma_{n}f_{H}(\lambda),\qquad A_{n}:=T_{n}(a_{n}),$
where $T_{n}(a_{n})$ denotes the $n\times n$ Toeplitz matrix generated by the symbol $a_{n}$. Assume $n\Delta_{n}\to\infty$ and let $b,c\in\{f_{H},\;\dot{f}_{H}\}$. Define the logarithmic-index function
$k(f_{H}):=0,\qquad k(\dot{f}_{H}):=1,$
which counts the number of logarithmic Fisher–Hartwig factors appearing in the symbol (hence governs the possible $\log n$ loss in the remainder). Then
@@
$|\dot{f}_{H}(\lambda)|\leq f_{H}(\lambda)\,\bigl(2|\log|\lambda||+C_{1}\bigr)+C_{2}\,|\lambda|^{-p}\leq C\,f_{H}(\lambda)\bigl(1+|\log|\lambda||\bigr).$
@@ 3.2 Central Limit Theorem for $S_{\sigma,n}$
$J_{2}(H,\sigma)$ $:=\int_{\mathbb{R}}w(x)\Bigl(C_{H}-2\ln|x|\Bigr)^{2}\,dx,$ (30)
where $C_{H}$ is defined in equation (32) and given by $C_{H}=\partial_{H}\ln\left(c_{H}\right)$. Here the constant
$c_{H}=\frac{\Gamma(2H+1)\sin(\pi H)}{2\pi}.$
@@
Now we can deal with the part of the parameter $\sigma$ of the score funtion.
@@
$\left(\,\frac{\sqrt{2}\gamma_{n}}{\sigma}\left\lVert C_{n}\right\rVert_{F}\right)^{2}=\frac{2\sigma^{4}\Delta_{n}^{2p}}{\sigma^{2}}\cdot\frac{n}{2\pi}\Delta_{n}^{1-2p}\Big(J_{0}+o(1)\Big)=\frac{\sigma^{2}n\Delta_{n}}{\pi}\,\Big(J_{0}+o(1)\Big)=\frac{\sigma^{2}}{\pi}\,\mathcal{T}_{n}\Big(J_{0}+o(1)\Big).$
@@ 3.3 The CLT for the Remainder of Score function of $H$
$\Delta_{n}\,\kappa_{n}(x)b_{n}(x)^{2}$ $=\Delta_{n}^{1-2p}w(x)\Bigl(4\ln^{2}(1/\Delta_{n})+4\ln(1/\Delta_{n})\bigl(C_{H}-2\ln|x|\bigr)\Bigr)$
$\quad+\Delta_{n}^{1-2p}w(x)\bigl(C_{H}-2\ln|x|\bigr)^{2}+o\!\big(\Delta_{n}^{1-2p}\ln^{2}(1/\Delta_{n})\big).$ (40)
@@
$R_{H,n}:=\frac{\gamma_{n}}{2}\Big(Z_{n}^{\top}D_{n}Z_{n}-\mathrm{tr}(D_{n})\Big),\qquad Z_{n}\sim\mathcal{N}(0,I_{n}),$
@@
$\gamma_{n}^{2}\left\lVert D_{n}\right\rVert_{F}^{2}$ $=\sigma^{4}\Delta_{n}^{2p}\cdot\frac{n}{2\pi}\Delta_{n}^{1-2p}\Bigl(4\ln^{2}(1/\Delta_{n})\,J_{0}+{4\ln(1/\Delta_{n})\,J_{1}+J_{2}}+o(\ln^{2}(1/\Delta_{n}))\Bigr)$
$=\frac{2\sigma^{4}}{\pi}\,\mathcal{T}_{n}\,\ln^{2}(1/\Delta_{n})\Bigl(J_{0}+o(1)\Bigr),$